# Genetic algorithm and edge-colorings of complete graphs with connected classes

### v1

@@ Abstract
In this paper, we explore the application of a particular genetic algorithm, the Rank Genetic Algorithm (Rank GA), to address a graph theory problem. The Rank GA was introduced to address the problem of graph coloring, exploring into the specialized field of Chromatic Graph Theory. We successfully improved several of the previously known lower bounds. To validate the effectiveness of these new bounds, we performed an analytical approximation, confirming their validity. The findings underscore the interdisciplinary impact of the use of genetic algorithms in theoretical mathematical research.
Keywords: Rank GA, Genetic Algorithms for theoretical problems, Graph coloring, pseudoachromatic index, Hadwiger number
@@ 1 Introduction
Genetic Algorithms (GAs) have positioned themselves as powerful tools, taking advantage of efficient strategies to tackle complex optimization problems prevalent in real-world applications across diverse industries, including transportation and manufacturing [9]. Although GAs are widely recognized for their effectiveness in addressing practical engineering challenges, their interdisciplinary potential extends to the realm of theoretical advancements in mathematical research.
@@
On the other hand, within the field of mathematics, particularly Graph Theory, we focus on graph coloring research, exploring the specialized field of Chromatic Graph Theory. Our specific attention is directed towards the connected-pseudoachromatic index, denoted as $\psi_{c}^{\prime}(K_{n})$ (see 3). Determining $\psi_{c}^{\prime}(K_{n})$ is difficult, and so far, there have been very few exact results for $\psi_{c}^{\prime}(K_{n})$ of complete graphs with connected classes. However, it has been possible to bound it using analytical methods [2].
@@
The structure of this paper is as follows. In Section 2, we provide an overview of previous work in which a GA has been successfully used to solve graph theory problems. In Section 3 we state the problem and include the necessary definitions to understand it. In Section 4, we describe the Rank GA and in Section 5 we explain its use in our problem. We present the obtained results in Section 6, while in Section 7 we give an approximation of the upper bound. Finally, Section 8 contains the conclusions.
@@ 2 Background
In recent decades, GAs have emerged as effective tools for addressing graph theory problems with practical applications. Specifically, various GAs have been developed to tackle the Graph Coloring Problem (GCP), a fundamental issue in graph theory. In the simplest instance of the GCP, the objective is to assign colors to the vertices of a graph in such a way that no two adjacent vertices share the same color. The goal is to find a valid coloring of the graph with the minimal number of colors. Given its NP-hard nature and numerous practical applications, the GCP stands out as the most extensively researched combinatorial optimization problem in both computer science and mathematics [7]. In [11], you can find the most significant work in which GAs have been used to address the GCP. In addition to GCP, GAs have also been applied to other graph problems, among the most recent works are: the Shortest Path Problem [6], the Traveling Salesman Problem [12], Maximum Matching in bipartite graphs [10], and the Maximum Edge-Disjoint Path Problem [9].
@@ 3 Problem statement
The maximum number of colors in a connected complete coloring of a complete graph $K_{n}$ of order $n$ that can be achieved is called the connected-pseudoachromatic index and it is denoted $\psi_{c}^{\prime}(K_{n})$. There are certain values of $n$ for which it is known, and for others, only the lower and upper bounds are known [1, 2]. The problem we address here, in particular, consists in improving the known bounds [2] for some values of $n$.
@@ 3.1 Edge-colorings of a complete graph
The connected-pseudoachromatic index $\psi^{\prime}_{c}(G)$ of a graph $G$ is the maximum number of colors $k$ for which a connected complete edge-coloring exists in the graph $G$ using $k$ colors, when $G$ is connected, otherwise, it is defined as the maximum index over the indexes of its connected components. When we color vertices and each chromatic class is connected, the parameter is also known as the Hadwiger number of $G$.
@@ 4.1.3 Mutation by Rank
5 The Rank GA to find complete graph colorings with connected classes
5.1 Solution representation
In this problem, the individual is the edge coloring of a graph. Edge coloring can be represented by a coloring matrix of size $n\times n$, where $n$ is the number of vertices of the graph. Given a complete graph $G=(V,E)$, where $V=\{v_{0},v_{1},\ldots,v_{n-1}\}$, the coloring matrix $C(G)$ of $G$ is defined as the upper triangle of a square matrix of size $n\times n$ defined by $C(G)=[a_{ij}]$, $i<j$, where $a_{ij}$ is the color code of the edge between the vertices $i$ and $j$. The diagonal of this matrix is not considered, as there are no loops, and since it is a complete graph, the information in the lower triangle of the matrix is the same as that in the upper triangle; therefore, it is also not considered.
We call numAvailableColors the number of available colors, and numEdges the number of edges of the graph $G$. Each individual has numEdges-1 color codes, one for each edge of $G$. Each position in the individual genome is always related to the same edge in $G$. Each edge color is encoded by an integer number between 0 and numberAvailableColors - 1.
To save memory space, we represent each individual $X$ as an array of integers that contains the information of the coloring matrix. Thus, the first elements of the array are the color codes of the edges of the first row of $C(G)$ which are from $a_{0,1}$ to $a_{(0,\text{numVertex}-1)}$, then there are the edges of the second row from $a_{1,2}$ to $a_{(1,\text{numVertex}-1)}$, and so on up to $a_{(\text{numVertex}-2,\text{numVertex}-1)}$.
5.2 The fitness function
To evaluate an individual $X$, we employ three metrics. First, we use the numPresentColors metric, which is the number of colors present in the individual. In this context, it is desirable for the value of numPresentColors to be as large as possible. Second, we consider the number of chromatic classes that do not share a common vertex: numNotCommomVertex. Here, it is preferred for the value of numNotCommomVertex to be as small as possible. Third, numNotConnected is the number of chromatic classes present in the individual that are not connected, which must also be as small as possible.
The fitness function to maximize the fitness of individual $X$ is defined by Equation (5.1) as follows:
[…]
@@ 6 Results
Figure 2 visually represents the values detailed in Table 2. As can be seen, the known lower bound presents sudden steps at certain graph sizes, while the improved lower bounds present smother growth.
@@
If the reader wants to see some example solutions, please go to Appendix A.
@@ 7 On the upper bound
If $n\geq 8$ then an approximate upper bound is
@@ 8 Conclusions
The problem at hand involves improving, by adapting a genetic algorithm called the Rank GA, the known bounds of the connected pseudo-achromatic index, denoted as $\psi_{c}^{\prime}(K_{n})$, which represents the maximum number of colors in a connected and complete coloring of a complete graph $K_{n}$ of order $n$. The Rank GA demonstrated significant effectiveness in improving $\psi_{c}^{\prime}(K_{n})$ bounds for complete graphs with connected classes. The algorithm was able to find the solutions due to its balance in exploring the solution space and exploiting promising solutions. The improved bounds of $\psi_{c}^{\prime}(K_{n})$ obtained by the Rank GA were validated against an analytical approximation, confirming their validity. The success of the Rank GA in enhancing the bounds of $\psi_{c}^{\prime}(K_{n})$ may significantly contribute to theoretical advancements in Chromatic Graph Theory. Upon examination of the resulting colorings, it is observed that the sizes of the chromatic classes vary, with most sizes equal to the minimum size stipulated by Theorem 7.1. However, their distribution does not align with the color patterns described in Theorem 7.2, meaning that they lack a projective plane structure in their arrangement.

### v2

@@ Abstract
In this study, the Rank Genetic Algorithm (Rank GA) was adapted to address a problem in the field of Chromatic Graph Theory: the connected pseudo-pseudoachromatic index. We successfully improved several previously known lower bounds of that index. To validate the effectiveness of these new bounds, we performed an analytical approximation, confirming their validity. The findings highlight the interdisciplinary impact of the use of genetic algorithms in theoretical mathematical research, particularly in the solution of complex combinatorial problems. Our work demonstrates that computational methods can solve problems that are intractable to humans, leading to scientific advances in many areas. Moreover, the problem we addressed showcases the robustness of the Rank GA in finding optimal solutions within a reasonable time frame, as it consistently avoids getting stuck in local optima.
Keywords: Genetic Algorithm, Rank GA, Chromatic Graph Theory, Edge-Coloring, Connected Classes
@@ 1 Introduction
Genetic Algorithms (GA) have emerged as powerful tools, taking advantage of efficient strategies to address complex optimization problems prevalent in real-world applications in diverse industries, including transportation and manufacturing [9]. Although GAs are widely recognized for their effectiveness in addressing practical engineering challenges, their interdisciplinary potential extends to the realm of theoretical advancements in mathematical research.
@@
On the other hand, within the field of mathematics, particularly Graph Theory, we focus on graph coloring research, exploring the specialized field of Chromatic Graph Theory. Our specific attention is directed towards the connected-pseudoachromatic index, denoted $\psi_{c}^{\prime}(K_{n})$. Determining $\psi_{c}^{\prime}(K_{n})$ is difficult, and so far, there have been very few exact results for $\psi_{c}^{\prime}(K_{n})$ of complete graphs with connected classes. However, it has been possible to bound it using analytical methods [2].
@@
The structure of this paper is as follows. In Section 2, we provide an overview of previous work in which a GA has been successfully used to solve graph theory problems. In Section 3, we state the problem and include the necessary definitions to understand it. In Section 4, we describe the Rank GA and in Section 5, we present the obtained results. Finally, Section 7 contains the conclusions.
@@ 2 Background
In recent decades, GAs have emerged as effective tools to address problems of graph theory with practical applications. Specifically, various GAs have been developed to tackle the Graph Coloring Problem (GCP), a fundamental issue in graph theory. In the simplest instance of the GCP, the objective is to assign colors to the vertices of a graph in such a way that no two adjacent vertices share the same color. The goal is to find a valid coloring of the graph with the minimal number of colors. Given its NP-hard nature and numerous practical applications, the GCP stands out as the most extensively researched combinatorial optimization problem in both computer science and mathematics [7]. In [12], you can find the most significant work in which GAs have been used to address the GCP. In addition to GCP, GAs have also been applied to other graph problems, among the most recent works are: the Shortest Path Problem [6], the Traveling Salesman Problem [13], Maximum Matching in bipartite graphs [11], and the Maximum Edge-Disjoint Path Problem [9].
@@ 3 Problem Statement
The maximum number of colors in a connected complete coloring of a complete graph $K_{n}$ of order $n$ that can be achieved is called the connected pseudo-achromatic index and is denoted $\psi_{c}^{\prime}(K_{n})$. There are certain values of $n$ for which it is known, and for others, only the lower and upper bounds are known [1, 2]. The problem we address here, in particular, consists in improving the known bounds [2] for some values of $n$.
@@ 3.1 Edge-Colorings of a Complete Graph
The connected-pseudoachromatic index $\psi^{\prime}_{c}(G)$ of a graph $G$ is the maximum number of colors $k$ for which a connected complete edge-coloring exists in the graph $G$ using $k$ colors, when $G$ is connected; otherwise, it is defined as the maximum index over its connected components. When we color vertices and each chromatic class is connected, the parameter is also known as the Hadwiger number of $G$.
@@ 5 Results
Figure 2 visually represents the values detailed in Table 2. As can be seen, the known lower bound presents sudden steps at certain graph sizes, while the improved lower bounds present smoother growth.
@@ 6 On the Upper Bound
If $n\geq 8$, then an approximate upper bound is
@@ 7 Conclusions
The problem at hand involves improving, by adapting a genetic algorithm called Rank GA, the known bounds of the connected pseudoachromatic index, denoted $\psi_{c}^{\prime}(K_{n})$, which represents the maximum number of colors in a connected and complete coloring of a complete graph $K_{n}$ of order $n$. The Rank GA demonstrated significant effectiveness in improving $\psi_{c}^{\prime}(K_{n})$ bounds for complete graphs with connected classes. The algorithm was able to find the solutions due to its balance in exploring the solution space and exploiting promising solutions. The improved bounds of $\psi_{c}^{\prime}(K_{n})$ obtained by the Rank GA were validated against an analytical approximation, confirming their validity. The success of the Rank GA in enhancing the bounds of $\psi_{c}^{\prime}(K_{n})$ may significantly contribute to theoretical advancements in Chromatic Graph Theory. Upon examination of the resulting colorings, it is observed that the sizes of the chromatic classes vary, with most sizes equal to the minimum size stipulated by Theorem 6.1. However, their distribution does not align with the color patterns described in Theorem 6.2, meaning that they lack a projective plane structure in their arrangement.